OpenStudy (anonymous):

The sum of three consecutive integers is 90. Find the three numbers by letting x represent the first integer

OpenStudy (texaschic101):

x = 1st integer x + 1 = 2nd integer x + 2 = 3rd integer Now it says the sum of the integers equals 90, so lets set up the equation... (x) + (x + 1) + (x + 2) = 90 now just combine like terms and solve for x. Once x is found, sub that answer back into the 2nd and 3rd integers to find them.